# Yau’s Affine Normal Descent: Algorithmic Framework and Convergence Analysis

### 30 Mar 2026 v1

@@ 2.1.2. Slice-centroid formula (geometric approximation)
$\widehat{d}_{\mathrm{SC}}(z)\ \propto\ \frac{z-g(-\delta)}{\delta},$
whose truncation error is $O(\delta)$. This formula yields the analytic affine normal (up to scale) precisely when $z$ is an elliptic point [^1] A point $z$ is called elliptic if the tangent–tangent Hessian at $z$ is positive definite; in this case $M$ is locally strictly convex. It is called hyperbolic if that block is indefinite, and parabolic if it is singular., i.e., when the slice $S_{C}$ is convex for $C<0$ and shrinks smoothly to $z$. Outside the elliptic region (hyperbolic or parabolic points), the slices may be nonconvex or multi-component, and $d_{\rm SC}(z)$ is no longer aligned with the analytic affine normal. Note that this method requires only first-order information (for constructing the tangent plane) and avoids third derivatives and matrix inversion, though computing the high-dimensional centroid $g(C)$ efficiently remains the main bottleneck.
@@ 2.2. Equivalence of the analytical and slice-centroid formulas under convexity
The slice-centroid formula captures the affine normal direction only when the level hypersurface is locally strictly convex. In general nonconvex situations, tangential drift may occur and the slice-centroid direction may fail to agree with the analytical affine-normal direction; see Subsection 1.3 for discussion.
@@
where $P(x)=\nabla f(z)\cdot(x-z)$ and $C<0$ corresponds to the interior. Then the inward one-sided limit
$\lim_{C\uparrow 0}\frac{z-g(C)}{-C}$
exists and agrees with the analytical affine-normal direction at $z$, up to a positive scalar multiple.
@@
Take the parallel slice planes $P_{C}=\{(u,t):\ t=C\}$ with small $C>0$. Then
$S_{C}=\{(u,C):\ \Phi(u)\leq C\}$
@@
$g(C)=(\bar{u}(C),C),\qquad\bar{u}(C)=\frac{1}{V(C)}\int_{\{\Phi(u)\leq C\}}u\,du,\quad V(C)=\int_{\{\Phi(u)\leq C\}}1\,du.$
Step 3 (Scaling). Let $u=\sqrt{C}\,y$ to obtain
$\Phi(\sqrt{C}\,y)=\tfrac{1}{2}C|y|^{2}+\tfrac{1}{6}C^{3/2}C_{ijk}y_{i}y_{j}y_{k}+O(C^{2}).$
Set $\varepsilon=\sqrt{C}$ and define
@@
$V(C)=C^{n/2}|\mathbb{D}_{\varepsilon}|,\qquad\int_{\{\Phi(u)\leq C\}}u\,du=C^{(n+1)/2}\int_{\mathbb{D}_{\varepsilon}}y\,dy,$
@@
$\bar{u}(C)=\sqrt{C}\,\frac{\int_{\mathbb{D}_{\varepsilon}}y\,dy}{|\mathbb{D}_{\varepsilon}|}.$
@@
$\bar{u}(C)=\sqrt{C}\,\frac{O(\varepsilon^{2})}{|\mathbb{B}|+O(\varepsilon^{2})}=O(C^{3/2}),$
@@
$\lim_{C\downarrow 0}\frac{z-g(C)}{-C}=(0,\ldots,0,1)\in\mathbb{R}^{n+1}.$
Step 7 (Recovering the affine normal). Under the equi-volume affine normalization, $(0,\ldots,0,1)$ is the analytical affine-normal direction. Undoing the normalization preserves direction up to scale. ∎
@@ 2.3. Role of convexity
Failure at non-elliptic points. At hyperbolic points, the tangent-tangent Hessian is indefinite: the level set bends in opposite directions, slices become nonconvex or multi-component, and their centroids no longer reflect the analytic affine normal. At parabolic points, the tangent-tangent Hessian is singular and the affine metric degenerates, so neither the analytic affine normal nor the moment-based construction yield a meaningful normal direction. The counterexample in Example 4.3 illustrates that the slice-centroid direction may even become an ascent direction in the nonconvex case.
Implications for descent directions in optimization. The analytic affine normal remains formally well defined at any nondegenerate point of the hypersurface, but it represents an inward geometric direction only at elliptic points. Consequently, the affine normal is a guaranteed strict descent direction for $f$ if each iterate $x_{k}$ lies on a locally strictly convex patch of the level set. Within such a neighborhood, the interior sublevel set $\{f\leq f(x_{k})\}$ lies on the inward side of the hypersurface, and the affine normal points strictly into this region. Moreover, in the same elliptic neighborhood, the slice-centroid and cap-centroid constructions provide consistent approximations of the analytic affine normal, since all three directions coincide up to scaling.
@@ 4.2. Moment-based constructions require convexity
Example 4.3 (Slice-centroid fails to be a descent direction in the nonconvex case).
@@
$f(x,y)=(x^{2}-1)^{2}+y-1,\qquad z=(0,0),$
for which $f(z)=0$ and $\nabla f(z)=(0,1)$. The tangent line at $z$ is $y=0$, and
$\Omega_{z}=\{(x,y):f(x,y)\leq 0\}=\{y\leq 1-(x^{2}-1)^{2}\}$
@@
$S_{C}=\{(x,C):(x^{2}-1)^{2}\leq 1-C\}$
consists of two disconnected symmetric intervals, hence its centroid is $g(C)=(0,C)$. Thus the slice-centroid direction
$\widehat{d}_{\mathrm{SC}}(z)\propto\frac{g(0)-g(-\delta)}{\delta}=(0,1)$
@@
$\langle\nabla f(z),\widehat{d}_{\mathrm{SC}}(z)\rangle=1>0,$
an ascent direction. This reflects the non-ellipticity at $z$ and shows that the slice-centroid construction require local convexity.
@@ 10.4.1. Rosenbrock function
For comparison, Figures 9 and 10 illustrate the behavior of gradient descent and damped Newton’s method under the same Wolfe conditions. Gradient descent exhibits the well-known zigzagging behavior when traversing the narrow valley, resulting in slow progress. Damped Newton accelerates once the iterates approach the minimizer, but requires substantial damping in the early phase to maintain stability.
@@ Acknowledgements
Y.-S. N. was supported by the National Natural Science Foundation of China (Grant No. 42450242) and the Beijing Overseas High-Level Talent Program. A. S. would like to acknowledge support from the Beijing Natural Science Foundation (Grant No. BJNSF–IS24005) and the NSFC–RFIS Program (Grant No. W2432008). He also thanks the NSF AI Institute for Artificial Intelligence and Fundamental Interactions at the Massachusetts Institute of Technology (MIT), funded by the U.S. National Science Foundation under Cooperative Agreement PHY–2019786, as well as China’s National Program of Overseas High-Level Talent for generous support. All three authors gratefully acknowledge institutional support from the Beijing Institute of Mathematical Sciences and Applications (BIMSA). The authors would also like to thank Prof. Liping Zhang of Tsinghua University for helpful discussions.

### 03 Apr 2026 v2

@@ 2.1.2. Slice-centroid formula (geometric approximation)
$\widehat{d}_{\mathrm{SC}}(z)\ \propto\ \frac{g(-\delta)-z}{\delta},$
whose truncation error is $O(\delta)$. This formula yields the analytic affine normal (up to scale) precisely when $z$ is an elliptic point [^1] A point $z$ is called elliptic if the tangent–tangent Hessian at $z$ is positive definite; in this case $M$ is locally strictly convex. It is called hyperbolic if that block is indefinite, and parabolic if it is singular., i.e., when the slice $S_{C}$ is convex for $C<0$ and shrinks smoothly to $z$. Outside the elliptic region (hyperbolic or parabolic points), the slices may be nonconvex, disconnected, or even unbounded, and $d_{\rm SC}(z)$ is no longer guaranteed to align with the analytic affine normal. Under the inward convention, whenever the centroid exists, the slice-centroid direction still has an inward/descent normal component, since
$\Bigl\langle\nabla f(z),\,\frac{g(C)-z}{-C}\Bigr\rangle=\frac{P(g(C))}{-C}=\frac{C}{-C}=-1.$
Thus the real difficulty outside the elliptic regime is geometric well-posedness and consistency, not the descent sign alone. Note that this method requires only first-order information (for constructing the tangent plane) and avoids third derivatives and matrix inversion, though computing the high-dimensional centroid $g(C)$ efficiently remains the main bottleneck.
@@ 2.2. Equivalence of the analytical and slice-centroid formulas under convexity
The slice-centroid formula captures the affine normal direction only when the level hypersurface is locally strictly convex. In general nonconvex situations, the slices may cease to be convex bodies: they can become disconnected, nonconvex, or unbounded, and the centroid trajectory may fail to agree with the analytical affine-normal direction. Under the inward convention, whenever the centroid exists, the corresponding slice-centroid direction is automatically inward in the normal component; the real issue is therefore geometric well-posedness and consistency, not descent sign alone.
@@
where $P(x)=\nabla f(z)\cdot(x-z)$ and $C<0$ corresponds to the interior. Then the inward one-sided limit defining $d_{\mathrm{SC}}(z)$ is well defined, and
$d_{\mathrm{SC}}(z)\ \parallel\ d_{\mathrm{AN}}(z).$
@@
Take the parallel slice planes $P_{r}=\{(u,t):\ t=r\}$ with small $r>0$. Then
$S_{r}=\{(u,r):\ \Phi(u)\leq r\}$
@@
$g(r)=(\bar{u}(r),r),\qquad\bar{u}(r)=\frac{1}{V(r)}\int_{\{\Phi(u)\leq r\}}u\,du,\quad V(r)=\int_{\{\Phi(u)\leq r\}}1\,du.$
Step 3 (Scaling). Let $u=\sqrt{r}\,y$ to obtain
$\Phi(\sqrt{r}\,y)=\tfrac{1}{2}r|y|^{2}+\tfrac{1}{6}r^{3/2}C_{ijk}y_{i}y_{j}y_{k}+O(r^{2}).$
Set $\varepsilon=\sqrt{r}$ and define
@@
$V(r)=r^{n/2}|\mathbb{D}_{\varepsilon}|,\qquad\int_{\{\Phi(u)\leq r\}}u\,du=r^{(n+1)/2}\int_{\mathbb{D}_{\varepsilon}}y\,dy,$
@@
$\bar{u}(r)=\sqrt{r}\,\frac{\int_{\mathbb{D}_{\varepsilon}}y\,dy}{|\mathbb{D}_{\varepsilon}|}.$
@@
$\bar{u}(r)=\sqrt{r}\,\frac{O(\varepsilon^{2})}{|\mathbb{B}|+O(\varepsilon^{2})}=O(r^{3/2}),$
@@
$\lim_{r\downarrow 0}\frac{g(r)-z}{r}=(0,\ldots,0,1)\in\mathbb{R}^{n+1}.$
Step 7 (Recovering the affine normal). Under the equi-volume affine normalization, $(0,\ldots,0,1)$ is the analytical affine-normal direction. Undoing the normalization preserves direction up to scale. Returning to the original notation C=r<0C=-r<0 for inward slices gives
$\lim_{C\uparrow 0}\frac{g(C)-z}{-C}\ \parallel\ d_{\mathrm{AN}}(z).$
@@ 2.3. Role of convexity
Failure at non-elliptic points. At hyperbolic points, the tangent-tangent Hessian is indefinite: the level set bends in opposite directions, slices become nonconvex, disconnected, or unbounded, and their centroids no longer reliably reflect the analytic affine normal. At parabolic points, the tangent-tangent Hessian is singular and the affine metric degenerates, so neither the analytic affine normal nor the moment-based construction yield a meaningful normal direction. Example 4.3 illustrates that in the nonconvex case the slice-centroid construction may remain descent in its normal component while still becoming geometrically unreliable because the slices cease to be convex bodies.
Implications for descent directions in optimization. The analytic affine normal remains formally well defined at any nondegenerate point of the hypersurface, but it represents an inward geometric direction only at elliptic points. Consequently, the analytic affine normal is a guaranteed strict descent direction for $f$ if each iterate $x_{k}$ lies on a locally strictly convex patch of the level set. By contrast, under the inward slice convention, the slice-centroid direction automatically carries a descent normal component whenever the centroid is well defined; however, outside the elliptic regime its slices may fail to be convex bodies and the resulting moment-based direction need not approximate the analytic affine normal. Within an elliptic neighborhood, the slice-centroid and cap-centroid constructions do provide consistent approximations of the analytic affine normal, since all three directions coincide up to scaling.
@@ 4.2. Moment-based constructions require convexity
Example 4.3 (Disconnected slices in the nonconvex case).
@@
$f(x,y)=x^{2}(x^{2}-1)(x^{2}-4)(x^{2}-9)-y,\qquad z=(0,0),$
for which $f(z)=0$ and $\nabla f(z)=(0,-1)$. The tangent line at $z$ is $y=0$, and
$\Omega_{z}=\{(x,y):f(x,y)\leq 0\}={\Bigl\{}(x,y):y\geq x^{2}(x^{2}-1)(x^{2}-4)(x^{2}-9){\Bigr\}}$
@@
$S_{C}={\Bigl\{}(x,-C):x^{2}(x^{2}-1)(x^{2}-4)(x^{2}-9)\leq-C{\Bigr\}}$
may already be a union of several disconnected symmetric intervals. For example, when C=24C=-24, the slice is disconnected, as shown in Figure 2. By symmetry, whenever the centroid exists it satisfies g(C)=(0,C)g(C)=(0,-C). Thus the slice-centroid direction
$\widehat{d}_{\mathrm{SC}}(z)\propto\frac{g(C)-z}{-C}=(0,1)$
@@
$\langle\nabla f(z),\widehat{d}_{\mathrm{SC}}(z)\rangle=-1<0,$
so under the inward convention it is indeed a descent direction. However, the slices are disconnected and not convex bodies. Thus the moment-based construction is no longer a faithful local geometric proxy for the analytic affine normal: the issue in the nonconvex case is not automatic loss of descent, but rather the loss of convexity, connectedness, and geometric well-posedness of the slices.
Figure 2. Disconnected inward slices for $f(x,y)=x^{2}(x^{2}-1)(x^{2}-4)(x^{2}-9)-y$ at $z=(0,0)$. The red curve is the zero level set $f(x,y)=f(z)$, and the green set is a representative slice $S_{C}$ for some $C<0$. Under the inward convention, the centroid direction remains inward/descent, but the slice is disconnected and thus ceases to be a faithful local geometric proxy for the analytic affine normal.
@@ 10.4.1. Rosenbrock function
For comparison, Figures 10 and 11 illustrate the behavior of gradient descent and damped Newton’s method under the same Wolfe conditions. Gradient descent exhibits the well-known zigzagging behavior when traversing the narrow valley, resulting in slow progress. Damped Newton accelerates once the iterates approach the minimizer, but requires substantial damping in the early phase to maintain stability.
@@ Acknowledgements
Y.-S. N. was supported by the National Natural Science Foundation of China (Grant No. 42450242) and China’s National Program of Overseas High-Level Talent. A. S. would like to acknowledge support from the Beijing Natural Science Foundation (Grant No. BJNSF–IS24005) and the NSFC–RFIS Program (Grant No. W2432008). He also thanks the NSF AI Institute for Artificial Intelligence and Fundamental Interactions at the Massachusetts Institute of Technology (MIT), funded by the U.S. National Science Foundation under Cooperative Agreement PHY–2019786, as well as China’s National Program of Overseas High-Level Talent for generous support. All three authors gratefully acknowledge institutional support from the Beijing Institute of Mathematical Sciences and Applications (BIMSA). The authors would also like to thank Prof. Liping Zhang of Tsinghua University for helpful discussions.